# Exploring Holomorphic Retracts

### v6

@@ Conclusion
As mentioned in the introduction until this stage, we have for the most part, restricted attention to retracts of bounded domains; though the domains in the previous section allows topological non-triviality, as we considered ‘analytic complements’ $D\setminus A$ at the beginning therein, the boundedness assumption prevailed. To go beyond, to more general settings, we require the following lemma:
@@
Let $Z$ be a retract of ×$\mathbb{C}\times D$ passing through $p:=(a,c)$. Let $F:\mathbb{C}\times D\to\mathbb{C}\times D$ be a holomorphic retraction map such that ${\rm image}(F)=Z$. Write $F(z,w)=\left(f_{1}(z,w),f_{2}(z,w)\right)$.
@@
Case-1: $g\equiv c$. In this case, $F(z,w)=(f_{1}(z,w),c)$. By comparing the first component of $F\circ F=F$, we get $f_{1}(f_{1}(z,w),c)=f_{1}(z,w)$. Note that $F$ is a holomorphic retraction map on ×$\mathbb{C}\times D$ onto $Z_{1}\times\{c\}$, where $Z_{1}:={\rm image}(f_{1})$ on ×$\mathbb{C}\times D$. By lemma 9.1, $Z_{1}$ is a retract of $\mathbb{C}$ passing through $a$. Hence either $Z_{1}=\mathbb{C}$, or $Z_{1}=\{a\}$. It follows that $Z=\mathbb{C}\times\{c\}$ or $Z=\{p\}$.
@@
Next we shall detail the complete determination of the retraction maps of ×$\mathbb{C}\times\Delta$.
@@
where $P_{i}(z,w)$ is a homogenous polynomial in $z,w$ of degree $i$.
@@
Next, equating degree $3$ terms on both sides, we get $P_{3}(P_{1},0)=0$. This implies that $P_{3}(z,w)=w(Q_{2}(z,w))$, where $Q_{2}(z,w)$ is a polynomial in $z$ and $w$. Continuing in this way, we get $P_{i}(z,w)=w(Q_{i-1}(z,w))$, where $Q_{i-1}$ is a polynomial in $z$ and $w$. Finally $f(z,w)=z+w(h(z,w))$, where $h$ is some holomorphic map on ×$\mathbb{C}\times\Delta$. In this subcase, $F(z,w)=(z+w(h(z,w)),0)$. Hence $Z=\mathbb{C}\times\{0\}$.
@@
where $P_{i}(z,w)$ is a homogenous polynomial in $z,w$ of degree $i$
@@
Since ×$\mathbb{C}\times\Delta$ is homogenous, upto a conjugation of automorphism, every non-trivial retraction map is of the form $F(z,w)=\left(z+wh(z,w),0\right)$, where $h\in\mathcal{O}(\mathbb{C}\times\Delta)$ or $F(z,w)=(h(w),w)$, where $h\in\mathcal{O}(\Delta)$
@@
Let $\widetilde{Z}$ be a retract of $(\mathbb{C}\setminus A)\times D$ passing through an arbitrary point $(a,b)\in(\mathbb{C}\setminus A)\times D$. Let $F:(\mathbb{C}\setminus A)\times D\to(\mathbb{C}\setminus A)\times D$ be a retraction map such that ${\rm image}(F)=\widetilde{Z}$. Write $F(z,w)=(f(z,w),g(z,w))$. For each fixed $w\in D$, $g_{w}:(\mathbb{C}\setminus A)\to D$ defined by $g_{w}(z)=g(z,w)$ is a bounded holomorphic map on $(\mathbb{C}\setminus A)$. Applying Riemann removable singularities theorem (for holomorphic functions of several complex variables) to each component of $g_{w}$, $g_{w}$ can extend holomorphically as a map $g_{w}^{\prime}:\mathbb{C}\to\overline{D}$. By Liouville’s theorem, $g_{w}^{\prime}$ is a constant function. In particular, $g_{w}$ is constant function and thus implies that $g(z,w)$ is independent of $z$. So $g(z,w)=\tilde{g}(w)$ for some holomorphic map $\tilde{g}$ on $D$. From the definition of retraction map, we get $F\circ F=F$. This implies in particular that $f\left(f(z,w),\tilde{g}(w)\right)=f(z,w)$ and $\tilde{g}(\tilde{g}(w))=\tilde{g}$. Since $F(a,b)=(a,b)$, we get $\tilde{g}(b)=b$. Note that $\tilde{g}$ is a retraction map on $D$ such that $\tilde{g}(b)=b$. For each fixed $w\in D$, we get a holomorphic map $f_{w}:(\mathbb{C}\setminus A)\to(\mathbb{C}\setminus A)$ defined by $f_{w}(z)=f(z,\tilde{g}(w))$. Note that
@@
This shows that $f_{w}$ is a retraction map on $(\mathbb{C}\setminus A)$. Since every domain in $\mathbb{C}$, the only retraction maps are identity and constant function, either $f_{w}\equiv I$, the identity map on $(\mathbb{C}\setminus A)$ or $f_{w}\equiv$ constant function (depends on $w$), hence imples that $f_{w}$ is independent of $z$. Hence for all $z,w\in(\mathbb{C}\setminus A)\times D$, we have
@@
Define $Z:={\rm image}(\tilde{g})$. According to (9.4) and (9.5), we divide into $2$ cases.
@@
We claim that $\widetilde{Z}=\{(H(w),w):w\in Z\}$. To prove this claim, observe that for each $w\in Z$, we have $\tilde{g}(w)=w$ and $f(H(w),w)=f(H(w),\tilde{g}(w))=H(w)$ This implies that $\{(H(w),w):w\in Z\}\subset\widetilde{Z}$. For attaining converse implication, let $(z_{0},w_{0})$ be an arbitrarily point in $\widetilde{Z}$. Since $\tilde{g}(w_{0})=w_{0}$ and $f(z_{0},w_{0})=z_{0}$ Hence we get $z_{0}=f(z_{0},w_{0})=f(z_{0},\tilde{g}(w_{0}))=H(\tilde{g}(w_{0}))=H(w_{0})$. This implies that $(z_{0},w_{0})=\left(H(w_{0}),w_{0}\right)$. Hence $\widetilde{Z}\subset\{\left(H(w),w\right):w\in Z\}$, and this proves the claim.
@@
We shall next record the proof of theorem 1.22, whose verifying details will be shortened as most of its arguments are contained in the prequel. To explicitly spell out the special case of domains contained in Euclidean space: taking $X$ to be the simplest non-compact Riemann surface namely $\mathbb{C}$, we have a description of all the retracts of ×$\mathbb{C}\times D$ where $D$ is any hyperbolic domain in $\mathbb{C}^{N}$. Indeed as mentioned earlier, this class contains domains with a wide variety of fundamental (or for that matter higher homotopy and homology) groups.
@@
We shall now detail the proof of theorem 1.18 about retracts of the Hartogs triangle and its analytic complements. Making the identification $\Omega_{H}\simeq\Delta\times\Delta^{*}$, retracts of the Hartogs triangle can be deduced as a special corollary of lemma 8.1 taking $D$ to be just the bidisc $\Delta^{2}$. However, for analytic complements $\Omega_{H}\setminus A$ this cannot be done merely by citing lemma 8.1 (at the very least, directly!), as the Hartogs triangle $\Omega_{H}$ does not admit a Stein neighbourhood basis. We therefore prefer to give a direct proof as follows.
@@
Let $F$ be a non-constant holomorphic retraction map on $(\Delta\times\Delta^{*})\setminus A$ such that $Z={\rm image}(F)$. Write $F(z,w)=\left(F_{1}(z,w),F_{2}(z,w)\right)$, where $F_{1}:(\Delta\times\Delta^{*})\setminus A\to\Delta$ and $F_{2}:(\Delta\times\Delta^{*})\setminus A\to\Delta^{*}$ are holomorphic functions. Since $A$ is a proper analytic subset of $\Delta\times\Delta^{*}$ and by Riemann removable singularities theorem, $F_{1},F_{2}$ can be extended holomorphically as a map $\widetilde{F}_{1},\widetilde{F}_{2}:\Delta\times\Delta^{*}\to\overline{\Delta}$ respectively. Write $\Delta\times\Delta^{*}=\Delta^{2}\setminus A_{0}$, where $A_{0}=\{(z,w)\in\Delta^{2}:w=0\}$. Note that $A_{0}$ is a proper analytic subset of $\Delta^{2}$. Applying the same argument, $\widetilde{F}_{1},\widetilde{F}_{2}$ can be extended to a holomorphic map from $\Delta^{2}$ to $\overline{\Delta}$. By abuse of notation, $\widetilde{F}_{1},\widetilde{F}_{2}$ are holomorphic maps from $\Delta^{2}$ to $\Delta$. We now observe that the images of these maps actually does not contain any point from $\partial\Delta$ i.e., they map into $\Delta$; this follows by the open mapping theorem (non-constant holomorphic functions from domains in $\mathbb{C}^{N}$ are open maps into $\mathbb{C}$;). Note that atleast one of $\widetilde{F}_{1}$ or $\widetilde{F}_{2}$ are non-constant! So, $\widetilde{F}_{1},\widetilde{F}_{2}$ are holomorpic maps from $\Delta^{2}$ to $\Delta$. Thereby, $F$ extends holomorphically as a map $\widetilde{F}:\Delta^{2}\to\Delta^{2}$; elaborately for the sequel, $\widetilde{F}(z,w)=\left(\widetilde{F}_{1}(z,w),\widetilde{F}_{2}(z,w)\right)$ to $\Delta^{2}$.
@@
Let $\widetilde{Z}:={\rm image}(\widetilde{F})$. By theorem 3 in [Heath_Suff], $\widetilde{Z}$ is either of the form $\{(z,f(z)):z\in\Delta\}$ for some holomorphic self map $f$ on $\Delta$ or $\{(g(w),w):w\in\Delta\}$ for some holomorphic self map $g$ on $\Delta$. Since $Z\subset(\Delta\times\Delta^{*})\setminus A$ and $Z\subset\widetilde{Z}$ ($\widetilde{F}$ is a holomorphic extension of $F$), we get $Z$ is either of the form $\{(z,f(z)):z\in\Delta\}\cap((\Delta\times\Delta^{*})\setminus A)$ for some holomorphic map $f:\Delta\to\Delta$ or $\{(g(w),w):w\in\Delta^{*}\}\cap((\Delta\times\Delta^{*})\setminus A)$ for some holomorphic map $g:\Delta\to\Delta$. ∎
@@
is a holomorphic retract of $\mathbb{C}^{2}$ that can neither be written as a graph of any holomorphic map on the first factor nor the second; indeed, it is the image of the holomorphic retraction map $f$ on $\mathbb{C}^{2}$ defined by $f(z,w)=(e^{zw},\;zwe^{-zw})$ and $Z$ can in-fact not be written as a graph of any holomorphic map on any one-dimensional linear subspace of $\mathbb{C}^{2}$.
@@
Let $L_{1}$ be an arbitrary ray emanating from origin and $z$ an arbitrary point of $\mathbb{C}\setminus L_{1}$. Just by the closedness of $L_{1}$ in $\mathbb{C}$, we may pick an open disc $D\subset\mathbb{C}\setminus L_{1}$. As $\mathbb{C}\setminus L_{1}$ is simply connected, there exists atleast one (indeed infinitely many) branch of the (complex) logarithm on $\mathbb{C}\setminus L_{1}$; pick one of them and call it $\ell_{1}$. Choose a ray $L_{2}$ starting from origin such that $L_{2}$ intersects $D$ Note that $L_{2}$ divides $D$ into two regions which shall be referred to by $D^{+}~{}\text{and}~{}D^{-}$. As with $\ell_{1}$, we likewise pick and denote by $\ell_{2}$, a branch of logarithm on $\mathbb{C}\setminus L_{2}$.
@@
We note that the above retract is however, straightenable and thereby conjugate to a graph – we say that a complex submanifold $M$ of $\mathbb{C}^{N}$ of dimension $k$ is straightenable or rectifiable if we can apply an automorphism of $\mathbb{C}^{N}$ which maps $M$ onto $\mathbb{C}^{k}$. Getting examples of retracts which are not conjugate to graphs is more non-trivial but it is not difficult to derive it from the work of Forstneric – Globevnik – Rosay [Forstn]. To briefly recall their main result, let us recall as they do in the introduction of [Forstn] that Abhyankar and Moh in [Abynkr] and M.Suzuki in [Suzuki] proved that if $P$ is a polynomial embedding of $\mathbb{C}$ into $\mathbb{C}^{2}$, then there exists a polynomial automorphism $\psi$ of $\mathbb{C}^{2}$ such that $\psi\circ P(\mathbb{C})=\mathbb{C}\times\{0\}$. This result is not true for holomorphic embeddings of $\mathbb{C}$. By proposition $1$ in [Forstn], there exists a proper holomorphic embedding $H:\mathbb{C}\to\mathbb{C}^{N}$ such that for no automorphism $\psi$ of $\mathbb{C}^{N}$, $\psi\circ H(\mathbb{C})=\mathbb{C}\times\{0\}\subset\mathbb{C}^{N}$. We want to prove that ${\rm image}(H)$ is a retract of $\mathbb{C}^{N}$. Towards this the manner in which this proper holomorphic embedding $H$ is defined and attained in the proof of proposition $2$ of their paper [Forstn] is now very briefly recalled. Start with the embedding $H_{0}:\mathbb{C}\to\mathbb{C}^{N}$ defined by $H_{0}(\zeta)=(\zeta,0)$. For $j\geq 1$, $H_{j}:=\psi_{j}\circ H_{j-1}$, where $\psi_{j}$ is some automorphism of $\mathbb{C}^{N}$. These $(H_{j})$’s consitute a sequence of proper embeddings of $\mathbb{C}$ satisfying some properties. Taking their uniform limit on compacts $H:=limH_{j}$ yields the desired embedding.
@@
Consider now, the holomorphic retraction $\rho_{0}:\mathbb{C}^{N}\to\mathbb{C}^{N}$ defined by $\rho_{0}(z,w)=(z,0)$, where $z\in\mathbb{C}$ and $w\in\mathbb{C}^{N-1}$. Note that ${\rm image}(H_{0})={\rm image}(\rho_{0})$. For $j\geq 1$, let $\rho_{j}:\mathbb{C}^{N}\to\mathbb{C}^{N}$ defined by $\rho_{j}:=\psi_{j}\circ H_{j}\circ\psi_{j}^{-1}$, where $\psi_{j}$ is the automorphism mentioned in the above construction. Observe that each $\rho_{j}$ is a holomorphic retraction with ${\rm image}(\rho_{j})={\rm image}(H_{j})$. Note that the $\rho_{j}$’s converge uniformly on compacts. If we denote the limit map by $\rho$ then, $\rho$ is also a retraction – this follows by passing to the limit in the defining equation of a retraction wherein taking a double limit is justified owing to the uniform convergence. Finally, we note that the limit of this retraction map $\rho$ satisfies ${\rm image}(H)={\rm image}(\rho)$. Hence we have verfied the existence of a one-dimensional holomorphic retract $Z$ of $\mathbb{C}^{N}$ which is not straightenable, thereby not conjugate to a graph as well.

### v7

@@ Conclusion
As mentioned in the introduction until this stage, we have for the most part, restricted attention to retracts of bounded domains; though the domains in the previous section allow topological non-triviality, as we considered ‘analytic complements’ $D\setminus A$ at the beginning therein, the boundedness assumption prevailed. To go beyond, to more general settings, we require the following lemma:
@@
Let $Z$ be a retract of $\mathbb{C}\times D$ passing through $p:=(a,c)$. Let $F:\mathbb{C}\times D\to\mathbb{C}\times D$ be a holomorphic retraction map such that ${\rm image}(F)=Z$. Write $F(z,w)=\left(f_{1}(z,w),f_{2}(z,w)\right)$.
@@
Case-1: $g\equiv c$. In this case, $F(z,w)=(f_{1}(z,w),c)$. By comparing the first component of $F\circ F=F$, we get $f_{1}(f_{1}(z,w),c)=f_{1}(z,w)$. Note that $F$ is a holomorphic retraction map on $\mathbb{C}\times D$ onto $Z_{1}\times\{c\}$, where $Z_{1}:={\rm image}(f_{1})$ on $\mathbb{C}\times D$. By lemma 9.1, $Z_{1}$ is a retract of $\mathbb{C}$ passing through $a$. Hence either $Z_{1}=\mathbb{C}$, or $Z_{1}=\{a\}$. It follows that $Z=\mathbb{C}\times\{c\}$ or $Z=\{p\}$.
@@
Next we shall detail the complete determination of the retraction maps of $\mathbb{C}\times\Delta$.
@@
where $P_{i}(z,w)$ is a homogeneous polynomial in $z,w$ of degree $i$.
@@
Next, equating degree $3$ terms on both sides, we get $P_{3}(P_{1},0)=0$. This implies that $P_{3}(z,w)=w(Q_{2}(z,w))$, where $Q_{2}(z,w)$ is a polynomial in $z$ and $w$. Continuing in this way, we get $P_{i}(z,w)=w(Q_{i-1}(z,w))$, where $Q_{i-1}$ is a polynomial in $z$ and $w$. Finally $f(z,w)=z+w(h(z,w))$, where $h$ is some holomorphic map on $\mathbb{C}\times\Delta$. In this subcase, $F(z,w)=(z+w(h(z,w)),0)$. Hence $Z=\mathbb{C}\times\{0\}$.
@@
where $P_{i}(z,w)$ is a homogeneous polynomial in $z,w$ of degree $i$
@@
Since $\mathbb{C}\times\Delta$ is homogeneous, upto a conjugation of automorphism, every non-trivial retraction map is of the form $F(z,w)=\left(z+wh(z,w),0\right)$, where $h\in\mathcal{O}(\mathbb{C}\times\Delta)$ or $F(z,w)=(h(w),w)$, where $h\in\mathcal{O}(\Delta)$
@@
Let $\widetilde{Z}$ be a retract of $(\mathbb{C}\setminus A)\times D$ passing through an arbitrary point $(a,b)\in(\mathbb{C}\setminus A)\times D$. Let $F:(\mathbb{C}\setminus A)\times D\to(\mathbb{C}\setminus A)\times D$ be a retraction map such that ${\rm image}(F)=\widetilde{Z}$. Write $F(z,w)=(f(z,w),g(z,w))$. For each fixed $w\in D$, $g_{w}:(\mathbb{C}\setminus A)\to D$ defined by $g_{w}(z)=g(z,w)$ is a bounded holomorphic map on $(\mathbb{C}\setminus A)$. Applying Riemann removable singularities theorem to each component of $g_{w}$, $g_{w}$ can extend holomorphically as a map $g_{w}^{\prime}:\mathbb{C}\to\overline{D}$. By Liouville’s theorem, $g_{w}^{\prime}$ is a constant function. In particular, $g_{w}$ is constant function and thus implies that $g(z,w)$ is independent of $z$. So $g(z,w)=\tilde{g}(w)$ for some holomorphic map $\tilde{g}$ on $D$. From the definition of retraction map, we get $F\circ F=F$. This implies in particular that $f\left(f(z,w),\tilde{g}(w)\right)=f(z,w)$ and $\tilde{g}(\tilde{g}(w))=\tilde{g}$. Since $F(a,b)=(a,b)$, we get $\tilde{g}(b)=b$. Note that $\tilde{g}$ is a retraction map on $D$ such that $\tilde{g}(b)=b$. For each fixed $w\in D$, we get a holomorphic map $f_{w}:(\mathbb{C}\setminus A)\to(\mathbb{C}\setminus A)$ defined by $f_{w}(z)=f(z,\tilde{g}(w))$. Note that
@@
This shows that $f_{w}$ is a retraction map on $(\mathbb{C}\setminus A)$. Since every domain in $\mathbb{C}$, the only retraction maps are identity and constant function, either $f_{w}\equiv I$, the identity map on $(\mathbb{C}\setminus A)$ or $f_{w}\equiv$ constant function (depends on $w$), hence implies that $f_{w}$ is independent of $z$. Hence for all $z,w\in(\mathbb{C}\setminus A)\times D$, we have
@@
[…]
@@
We claim that $\widetilde{Z}=\{(H(w),w):w\in Z\}$. To prove this claim, observe that for each $w\in Z$, we have $\tilde{g}(w)=w$ and $f(H(w),w)=f(H(w),\tilde{g}(w))=H(w)$ . This implies that $\{(H(w),w):w\in Z\}\subset\widetilde{Z}$. For attaining converse implication, let $(z_{0},w_{0})$ be an arbitrarily point in $\widetilde{Z}$. Since $\tilde{g}(w_{0})=w_{0}$ and $f(z_{0},w_{0})=z_{0}$ . Hence we get $z_{0}=f(z_{0},w_{0})=f(z_{0},\tilde{g}(w_{0}))=H(\tilde{g}(w_{0}))=H(w_{0})$. This implies that $(z_{0},w_{0})=\left(H(w_{0}),w_{0}\right)$. Hence $\widetilde{Z}\subset\{\left(H(w),w\right):w\in Z\}$, and this proves the claim.
@@
We shall next record the proof of theorem 1.24, whose verifying details will be shortened as most of its arguments are contained in the prequel. To explicitly spell out the special case of domains contained in Euclidean space: taking $X$ to be the simplest non-compact Riemann surface namely $\mathbb{C}$, we have a description of all the retracts of $\mathbb{C}\times D$ where $D$ is any hyperbolic domain in $\mathbb{C}^{N}$. Indeed as mentioned earlier, this class contains domains with a wide variety of fundamental (or for that matter higher homotopy and homology) groups.
@@
We shall now detail the proof of theorem 1.20 about retracts of the Hartogs triangle and its analytic complements. Firstly, a very quick proof is as follows. Making the identification $\Omega_{H}\simeq\Delta\times\Delta^{*}$, retracts of the Hartogs triangle can be deduced as an immediate corollary of lemma 8.1 taking $D$ to be just the bidisc $\Delta^{2}$. An alternative direct albeit longer proof is as follows.
@@
Let $F$ be a non-constant holomorphic retraction map on $(\Delta\times\Delta^{*})\setminus A$ such that $Z={\rm image}(F)$. Write $F(z,w)=\left(F_{1}(z,w),F_{2}(z,w)\right)$, where $F_{1}:(\Delta\times\Delta^{*})\setminus A\to\Delta$ and $F_{2}:(\Delta\times\Delta^{*})\setminus A\to\Delta^{*}$ are holomorphic functions. Since $A$ is a proper analytic subset of $\Delta\times\Delta^{*}$ and by Riemann removable singularities theorem, $F_{1},F_{2}$ can be extended holomorphically as a map $\widetilde{F}_{1},\widetilde{F}_{2}:\Delta\times\Delta^{*}\to\overline{\Delta}$ respectively. Write $\Delta\times\Delta^{*}=\Delta^{2}\setminus A_{0}$, where $A_{0}=\{(z,w)\in\Delta^{2}:w=0\}$. Note that $A_{0}$ is a proper analytic subset of $\Delta^{2}$. Applying the same argument, $\widetilde{F}_{1},\widetilde{F}_{2}$ can be extended to a holomorphic map from $\Delta^{2}$ to $\overline{\Delta}$. By a little abuse of notation, we view $\widetilde{F}_{1},\widetilde{F}_{2}$ are holomorphic maps from $\Delta^{2}$ to $\Delta$. We now observe that the images of these maps actually does not contain any point from $\partial\Delta$ i.e., they map into $\Delta$; this follows by the open mapping theorem (non-constant holomorphic functions from domains in $\mathbb{C}^{N}$ are open maps into $\mathbb{C}$;). Note that atleast one of $\widetilde{F}_{1}$ or $\widetilde{F}_{2}$ are non-constant! So, $\widetilde{F}_{1},\widetilde{F}_{2}$ are holomorpic maps from $\Delta^{2}$ to $\Delta$. Thereby, $F$ extends holomorphically as a map $\widetilde{F}:\Delta^{2}\to\Delta^{2}$; elaborately for the sequel, $\widetilde{F}(z,w)=\left(\widetilde{F}_{1}(z,w),\widetilde{F}_{2}(z,w)\right)$ to $\Delta^{2}$.
@@
[…]
@@
is a holomorphic retract of $\mathbb{C}^{2}$ that can neither be written as a graph of any holomorphic map on the first factor nor the second; indeed, it is the image of the holomorphic retraction map $f$ on $\mathbb{C}^{2}$ defined by $f(z,w)=(e^{zw},zw~e^{-zw})$ and $Z$ can in-fact not be written as a graph of any holomorphic map on any one-dimensional linear subspace of $\mathbb{C}^{2}$.
@@
Let $L_{1}$ be an arbitrary ray emanating from origin and $z$ an arbitrary point of $\mathbb{C}\setminus L_{1}$. Just by the closedness of $L_{1}$ in $\mathbb{C}$, we may pick an open disc $D\subset\mathbb{C}\setminus L_{1}$. As $\mathbb{C}\setminus L_{1}$ is simply connected, there exists atleast one (indeed infinitely many) branch of the (complex) logarithm on $\mathbb{C}\setminus L_{1}$; pick one of them and call it $\ell_{1}$. Choose a ray $L_{2}$ starting from origin such that $L_{2}$ intersects $D$ . Note that $L_{2}$ divides $D$ into two regions which shall be referred to by $D^{+}~\text{and}~D^{-}$. As with $\ell_{1}$, we likewise pick and denote by $\ell_{2}$, a branch of logarithm on $\mathbb{C}\setminus L_{2}$.
@@
We note that the above retract is however, straightenable and thereby conjugate to a graph – we say that a complex submanifold $M$ of $\mathbb{C}^{N}$ of dimension $k$ is straightenable or rectifiable if we can apply an automorphism of $\mathbb{C}^{N}$ which maps $M$ onto $\mathbb{C}^{k}$. Getting examples of retracts which are not conjugate to graphs is more non-trivial but it is not difficult to derive it from the work of Forstneric – Globevnik – Rosay [Forstn]. To briefly recall their main result, let us recall as they do in the introduction of [Forstn], that Abhyankar and Moh in [Abynkr] and M.Suzuki in [Suzuki] proved the following: if $P$ is a polynomial embedding of $\mathbb{C}$ into $\mathbb{C}^{2}$, then there exists a polynomial automorphism $\psi$ of $\mathbb{C}^{2}$ such that $\psi\circ P(\mathbb{C})=\mathbb{C}\times\{0\}$. This result is not true for holomorphic embeddings of $\mathbb{C}$ in $\mathbb{C}^{2}$ or for that matter in $\mathbb{C}^{N}$ for any $N\geq 2$. By the first proposition in [Forstn], there exists a proper holomorphic embedding $H:\mathbb{C}\to\mathbb{C}^{N}$ such that for no automorphism $\psi$ of $\mathbb{C}^{N}$, we can attain: $\psi\circ H(\mathbb{C})=\mathbb{C}\times\{0\}\subset\mathbb{C}^{N})$. For the embedding provided by that proposition, we prove here that ${\rm image}(H)$ is indeed a retract of $\mathbb{C}^{N}$. Towards this, the manner in which this proper holomorphic embedding $H$ is defined and attained in the proof of proposition $2$ of their paper [Forstn], is now very briefly recalled. Start with the standard embedding $H_{0}:\mathbb{C}\to\mathbb{C}^{N}$ defined by $H_{0}(\zeta)=(\zeta,0)$. For $j\geq 1$, let $H_{j}:=\psi_{j}\circ H_{j-1}$, where $\psi_{j}$ is some automorphism of $\mathbb{C}^{N}$. These $(H_{j})$’s constitute a sequence of proper embeddings of $\mathbb{C}$ satisfying some properties which we do not need to repeat here. Taking their uniform limit on compacts $H:=\lim H_{j}$ yields the desired embedding; for further details, we refer the reader to [Forstn] but this recap will suffice for our aforementioned purpose which we finish off in the next paragraph.
@@
Denote by $\rho_{0}:\mathbb{C}^{N}\to\mathbb{C}^{N}$, the simplest of standard one-dimensional retraction maps namely, $\rho_{0}(z,w)=(z,0)$, where $z\in\mathbb{C}$ and $w\in\mathbb{C}^{N-1}$. Note that ${\rm image}(H_{0})={\rm image}(\rho_{0})$. For $j\geq 1$, let $\rho_{j}:\mathbb{C}^{N}\to\mathbb{C}^{N}$ be defined by $\rho_{j}:=\psi_{j}\circ H_{j}\circ\psi_{j}^{-1}$, where $\psi_{j}$ is the automorphism mentioned in the above construction. Observe that each $\rho_{j}$ is a holomorphic retraction with ${\rm image}(\rho_{j})={\rm image}(H_{j})$. Note that the $\rho_{j}$’s converge uniformly on compacts. If we denote the limit map by $\rho$ then, $\rho$ is also a retraction – this follows by passing to the limit in the defining equation of a retraction for our specific map $\rho_{j}$ (i.e., writing $\rho_{j}\circ\rho_{j}=\rho_{j}$ elaborately), wherein taking a ‘triple limit’ is justified owing to the uniform convergence. Finally, we note that the limit of this retraction map $\rho$ satisfies ${\rm image}(H)={\rm image}(\rho)$. Hence we have verfied the existence of a one-dimensional holomorphic retract $Z$ of $\mathbb{C}^{N}$ which is not straightenable, thereby not conjugate to a graph as well.